# One-Shot Imitation under Mismatched Execution

### v2

@@ IV Approach
// Find closest snippet from play dataset
${{\hat{z}_{H}}}\leftarrow\operatorname*{arg\,min}\limits_{{\hat{z}_{H}}=\phi(v_{H}),v_{H}\in D_{\rm play}}{d({\hat{z}_{H}},{z_{R}^{i:i+K}}})$
// Append snippet to imagined demo
2      $\mathbf{\hat{z}_{H}}\text{.extend}({\hat{z}_{H}})$
Return Imagined demo embeddings $\mathbf{\hat{z}_{H}}$
Algorithm 1 Imagine-Demo: Retrieve and compose demonstration snippets given a robot trajectory
Input: Robot dataset ${D}_{\rm robot}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$
Output: Trained robot policy $\pi_{\theta}(a|s,\mathbf{z})$
1Initialize robot policy $\pi_{\theta}$ and paired dataset $D_{\rm paired}\leftarrow\{\}$ for $(\mathbf{v_{R}},\mathbf{\xi_{R}})\text{ in }D_{\rm robot}$ do
Generate robot embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$
// Retrieve and pair demonstration
// Add paired data into dataset
2      $D_{\rm paired}\leftarrow D_{\rm paired}\cup(\mathbf{\hat{z}_{H}},\mathbf{{z}_{R}},\mathbf{\xi_{R}})$
3while not converged do
// Get demo, robot embeddings with actions
$(\;\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}}\;)\sim D_{\rm paired}$
// Hybrid training
4      for $(s_{t},a_{t})$ in $\mathbf{\xi_{R}}$ do
// Condition on imagined demo
$\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{\hat{z}_{H}}))$
// Condition on robot video
5            $\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{z_{R}}))$
6
Return Trained Robot Policy $\pi$
Algorithm 2 RHyME: Train robot policy on both robot trajectories and imagined demonstrations
IV-A Training the Vision Encoder
We align the robot and demonstrator video embeddings in three ways: visually, temporally, and at the task level, all without requiring trajectory-level correspondences. We employ unsupervised losses $\mathcal{L}_{\rm vis}(\phi)$ and $\mathcal{L}_{\rm temp}(\phi)$ for visual and temporal alignment, following prior works [44, 45], and introduce an optional task alignment loss $\mathcal{L}_{\rm task}(\phi)$.
Figure 2: Cross-Embodiment Datasets (Left): We introduce three cross-embodiment demonstrators in the Franka Kitchen simulator with varying visual and execution mismatches. Sphere-Easy replaces the robot model with a sphere at the end effector location, inducing a visual gap. Sphere-Medium additionally deviates in manipulation style by moving objects along different trajectories. Sphere-Hard resembles humans, as the demonstrator is bimanual and parallelizes tasks. Results (Right): We report performance across all datasets on a range of baselines, showing that as the demonstrator’s execution visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperforms XSkill.
Visual Alignment. To align robot and demonstrator embeddings ($\mathbf{z_{R}}$, $\mathbf{z_{H}}$), we use SwAV [44], a self-supervised method that clusters images based on shared visual features. SwAV learns a set of $K$ prototype vectors, to which each image is assigned. The SwAV loss $\mathcal{L}_{\rm vis}(\phi)$ updates both the encoder and prototypes, aligning robot and demonstrator videos by clustering similar visual features.
Temporal Alignment. To align temporally adjacent frames in robot and demonstrator videos, we use Time Contrastive Loss [45]. This loss encourages embeddings of frames close in time to be similar. For each frame $z^{t}$, we define a positive set $\mathbf{z^{+}}$ of frames within a temporal window $w$, and a negative set $\mathbf{z^{-}}$ for frames outside this window. Using the contrastive loss $\mathcal{L}_{\rm temp}(\phi)$, we pull embeddings from the positive set closer and push negative set embeddings further apart, capturing temporal continuity across videos.
Task Alignment. Task-level alignment $\mathcal{L}_{\rm task}(\phi)$ is used when a small set of paired robot and demonstrator snippets is available. Unlike frame-level methods, this aligns video embeddings of the robot $\mathbf{z_{R}}$ and demonstrator $\mathbf{z_{H}}$. We compute the optimal transport distance $d(\mathbf{z_{R}},\mathbf{z_{H}})$ to measure the similarity between two sequences of video embeddings. We then apply a contrastive learning objective (INFO-NCE [46]) to pull matched embeddings closer and push different-task embeddings apart. The final task alignment loss is: $\mathcal{L}_{\rm task}(\phi)=-\sum\limits_{i}\frac{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))}{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))+\sum_{j\neq i}\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{j}))}$
Our final loss function for training the visual encoder $\phi$ is:
$\mathcal{L}(\phi)=\lambda_{\rm vis}\mathcal{L}_{\rm vis}(\phi)+\lambda_{\rm temp}\mathcal{L}_{\rm temp}(\phi)+\lambda_{\rm task}\mathcal{L}_{\rm task}(\phi)$ (1)
where $\lambda_{\rm task}=0$ by default and non-zero only with access to short-horizon paired data (Fig. 5).
IV-B Training the Robot Policy
Training Overview. Algorithm 2 provides an overview of our approach to train a robot policy $\pi_{\theta}$ using both robot trajectories and imagined demonstration videos. The training process has two stages.
Stage 1: Create a Paired Dataset. For each robot trajectory $\mathbf{\xi_{R}}$ and video $\mathbf{v_{R}}$ in $D_{\rm robot}$, we encode the robot video into embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$ using the learned video encoder $\phi$. We then retrieve imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ by aligning $\mathbf{z_{R}}$ with demonstration snippets from the play dataset $D_{\rm play}$, through the function Imagine-Demo. This produces a paired dataset $D_{\rm paired}$ containing $(\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}})$.
Stage 2: Train Policy on Paired Dataset. The policy $\pi_{\theta}$ is trained on the paired dataset $D_{\rm paired}$ in a hybrid fashion. For each element in $D_{\rm paired}$, we update the policy in two modes —Mode 1: The policy is conditioned on the robot video embeddings $\mathbf{z_{R}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{z_{R}})$, Mode 2: The policy is conditioned on the imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{\hat{z}_{H}})$. By alternating between these two modes, the policy learns to generalize from both robot and imagined demonstration videos, enabling it to handle execution mismatches.
Figure 3: Cross-Embodiment Vision Embeddings (Left): We use t-SNE to visualize cross-embodiment latent embeddings generated by robot and demonstrator when executing different tasks on all three datasets. TCC Failure Example (Right): The robot and video clip are equivalent, but specific frames have high TCC losses. For example, a frame showing the robot performing the ’kettle’ action has a high loss due to its nearest neighbor in the video performing both ’kettle’ and ’light’ actions. This frame cycles back to the robot performing ’light’, which is mismatched.
@@ IV-B Training the Robot Policy
Imagining Demonstration Videos. Algorithm 1 details the retrieval process for imaging a sequence of demonstration embeddings. We break the robot’s video into short-horizon windows and compare the embeddings with those from the play dataset, retrieving snippets with the lowest sequence-level distance. These retrieved snippets are concatenated to form an imagined long-horizon demonstration video. The key challenge is defining a distance function $d(\mathbf{z_{R}},\mathbf{z_{H}})$ that can handle video sequences of varying lengths. We propose two methods to compute this distance.
@@
Method 1: Optimal Transport Distance. We calculate the Wasserstein distance (Optimal Transport) [47] between the robot and demonstrator video embeddings, i.e., the cost of the optimal transport plan that transfers one sequence of video embeddings into another. The robot’s embedding distribution is defined as $\rho_{R}=\{1/T,1/T,\dots,1/T\}$, and the demonstrator’s embedding distribution is defined as $\rho_{H}=\{1/T^{\prime},1/T^{\prime},\dots,1/T^{\prime}\}$, where $T$ and $T^{\prime}$ are the lengths of the video sequences respectively. The cost function for the transport is $C\in\mathbb{R}^{T\times T^{\prime}}$ where $C^{ij}$ is the cosine distance between the robot embedding $z_{R}^{i}$ and the demonstration embedding $z_{H}^{j}$. Our goal is to find the optimal assignment $M\in\mathbb{R}^{T\times T^{\prime}}$ that transports the distribution from $\rho_{R}$ to $\rho_{H}$ while minimizing the cost of the plan. Formally, we need to find $M^{*}=\operatorname*{arg\,min}\limits_{M}\sum_{i}\sum_{j}C^{ij}M^{ij}$. After solving the optimal transport assignment, the distance function is the cost of the plan, i.e., $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{i}\sum_{j}C^{ij}M^{*ij}$. In practice, we optimize an entropy-regularized version of this problem to find an approximate solution efficiently using the Sinkhorn-Knopp algorithm [47].
@@
Method 2: Temporal-Cyclic Consistency (TCC) Distance. We calculate the TCC loss between human and robot videos following  [48] which computes cycle consistency between robot video embedding $\mathbf{z_{R}}=\{z_{R}^{1},z_{R}^{2},\dots,z_{R}^{T}\}$ and demonstrator $\mathbf{z_{H}}=\{z_{H}^{1},z_{H}^{2},\dots,z_{H}^{T^{\prime}}\}$. For each robot frame $z_{R}^{t}$, we first compute a similarity distribution $\mathbf{\alpha}$ of $z^{R}_{t}$ with respect to the demonstrator’s embeddings, to find a soft-nearest neighbor $\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}$. Then, $\tilde{z}_{H}$ cycles back to the robot video by again computing its similarity distribution $\mathbf{\beta}$ with respect to robot video embeddings to get its soft-nearest neighbor $\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$. The TCC distance for a robot frame $z_{R}^{t}$ is the mean square error with its cycled-back frame $\tilde{z}_{R}^{t}$ as $\ell_{\rm tcc}=||z_{R}^{t}-\tilde{z}_{R}^{t}||_{2}$. We define the video-level TCC distance function by summing over the frame-level losses $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{t=1}^{T}\ell_{\rm tcc}(z_{R}^{t})$.
@@
We hypothesize that video retrieval using TCC distance can be inaccurate in two cases: (1) When robot and demonstrator embeddings differ due to variations in execution speed or style, leading to poor frame alignment. (2) When multiple robot embeddings correspond to a single demonstrator frame, as in sequential robot tasks versus parallel demonstrator actions, causing ambiguity in matching.
@@ V Experiments
Simulation Setup and Datasets. We evaluate our approach using the Franka Kitchen simulator [49], where a 7-DOF Franka arm performs $7$ different tasks. We generate $3$ cross-embodiment video datasets each progressively increasing the embodiment and execution mismatch, which contain $580$ long-horizon ($\sim$25 seconds) robot trajectories completing a sequence of 4 tasks and a bank of demonstrator play data ($>$ 3 hours) for training our models. First, in Sphere-Easy, we replace the robot’s visual rendering with a sphere following the gripper’s position, creating a visual gap between robot and demonstrator. Second, in Sphere-Medium, we introduce manipulation style mismatches by applying randomized motion primitives to the demonstrator, such as the robot dragging an object while the demonstrator lifts and carries it. Finally, in Sphere-Hard, we create a further divergence where the demonstrator perform two tasks simultaneously, similar to how humans use two hands. Fig. 2 (left) illustrates the cross-embodiment datasets. Before training, all methods pre-train visual representations on unpaired robot and demonstrator videos, and policies are trained using Diffusion Policy[50].
@@
Baselines. XSkill [4] simply conditions on robot videos during train-time, and uses its shared representation space to generalize zero-shot to inputs of demonstrator videos at test-time. OraclePairing [7, 8] is the gold-standard approach, assuming an oracle pairs demonstration with a robot trajectory, enabling conditioning on the demonstrator at train time. Our approach, RHyME, finds a middle ground. Without pairing, it imagines demonstrator trajectories that perform the same tasks as a robot trajectory (Section IV-B) by exploiting sequence-level correspondences. We compare two variants of our algorithm, RHyME-TCC and RHyME-OT, which differ in their distance functions used for retrieval. We also show how vision representations can be improved using short-horizon robot-demonstrator task pairs.
@@
Evaluation and Success Metrics. At test-time, we provide the robot policy with demonstrator videos consisting of composition of 4 tasks not seen during training. We evaluate one-shot imitation performance across 20 different demonstrator videos, rolling out the robot policy from the last 5 model checkpoints, yielding 100 total trials per dataset. We measure Task Recall, which measures recall by counting the successfully completed tasks shown in the demonstrator’s video, and Task Imprecision, which measures imprecision and reports the percentage of tasks the robot attempts incorrectly—those not specified in the demonstrator’s video.
Q1. How does performance vary across different levels of execution mismatch?
As the demonstrator’s actions visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperform XSkill (Fig. 2). The OraclePairing baseline consistently achieves greater than $90\%$ task recall with minimal task imprecision. In the prescence of merely a cosmetic gap in Sphere-Easy, all methods have high task recall while maintaining precision. Still, we see that RHyME-OT has higher $90\%$ task recall compared to XSkill’s $82\%$. This gap increases on the Sphere-Medium dataset ($72\%$ vs $42\%$) and becomes larger on Sphere-Hard ($53\%$ vs $1\%$).
@@
Fig. 3 (left) investigates this trend by probing the visual representations of the video encoder $\phi$, common across policies. We plot the image embeddings of the robot and demonstrator across three different tasks using a t-SNE plot. In Sphere-Easy, robot and demonstrator embeddings are closely clustered by task execution, so the XSkill policy can zero-shot transfer to demonstration embeddings at test time, but quickly drops in performance as the execution mismatch increases. RHyME algorithms overcome this problem and successfully retrieve the correct demonstration videos at train-time by reasoning over sequences of embeddings. The demonstrator performs two tasks simultaneously in Sphere-Hard. Consequently, demonstrator embeddings during two simultaneous tasks lie between the individual task embedding clusters of the robot. RHyME-OT maintains its performance relative to baseline as a result, but we observe that RHyME-TCC suffers a significant drop in task recall compared to RHyME-OT ($34\%$ vs $53\%$). More interestingly, this method has much higher task imprecision, even in the easier Sphere-Medium dataset. This suggests inaccurate retrieval of demonstration videos at train-time.
@@
Figure 4: Optimal Transport Distances. We measure the similarity between robot and demonstrator videos on the Sphere-Hard dataset by computing the cost of the Optimal Transport (OT) plans. The sum over the entire transport cost matrix costs yields the distance between videos. OT costs are lowest when tasks are the same between videos (highlighted by a tick mark).
Q2. How does video retrieval using Optimal Transport and TCC impact policies at test-time?
When task embeddings are well clustered (Sphere-Easy, Fig. 3), TCC correctly cycles back and forth within clusters to yield low distance for videos matching in task specification, and high distance otherwise. However, as task embedding clusters deviate due to execution mismatches, we observe inaccuracies: in Sphere-Hard (Fig. 3 (right)) when both clips complete the same two tasks, a bimanual task embedding lies in between two robot task clusters which, resulting in cycling-back to the incorrect robot frame. We see that RHyME-TCC has relatively high task recall with lower task imprecision (due to incorrect retrievals) relative to other methods. RHyME-OT performs strictly better across datasets (Fig. 2).
@@
The key reason for this performance difference is that optimal transport computes distances by matching videos across a sequence of embeddings. Fig. 4 visualizes the cost of the optimal transport plan between prompt robot clips and demonstration videos in the hard Sphere-Hard dataset. Comparing a robot clip doing two tasks (e.g. kettle and light), the transport cost across assignments is minimum only when compared to the demonstrator performing those same two tasks. TCC, on the other hand, attempts to establish one-to-one correspondences between the robot and demonstration frames, which are lacking in this dataset.
@@
Q3. Does fine-tuning visual representations with task-equivalent pairs improve one-shot imitation?
Following Section IV-A, we perform a study assuming access to short-horizon task pairings across embodiments to fine-tune our representation space with the contrastive optimal transport-based task alignment loss $L_{task}(\phi)$ in the Sphere-Hard setting. We find that encouraging induced distributions over embeddings to be similar lifts the performance of both XSkill and RHyME-OT (Fig. 5), and scales up as more data is accessible for fine-tuning since task alignment helps accurately retrieve video snippets for policy training. Ultimately, comparing induced distributions over embeddings with optimal transport is a beneficial design choice for matching clips at a task-level in the face of execution mismatches, as RHyME-OT (0% fine-tuned) still significantly outperforms XSkill (fine-tuned).

### v5

@@ IV-B Training the Robot Policy
// Find closest short-horizon clip embedding in play dataset
$\mathbf{\mathbf{\hat{z}_{H}}}\leftarrow\operatorname*{arg\,min}_{\mathbf{z_{H}}\in D_{\rm play}}{d(\mathbf{z_{H}},{z_{R}^{i:i+K}}})$
// Extend imagined embedding sequence with retrieved demo
2      $\mathbf{\hat{z}_{H}}\text{.extend}(\mathbf{\hat{z}_{play}})$
Return Imagined Demo $\mathbf{\hat{z}_{H}}$
Algorithm 2 Imagine-Demo: Retrieving Matched Human Embeddings
@@
Imagining Human Demonstration Videos. Algorithm 2 details the retrieval process for imagining a sequence of human embeddings. We break the robot’s video into short-horizon windows and compare the embeddings with those from the play dataset, retrieving snippets with the lowest sequence-level distance. These retrieved snippets are concatenated to form an imagined long-horizon human demonstration video. The key challenge is defining a distance function $d(\mathbf{z_{R}},\mathbf{z_{H}})$ that can handle video sequences of varying lengths. We propose two methods to compute this distance: Optimal Transport Distance and Temporal Cyclic Consistency (TCC) Distance.
@@
Method 1: Optimal Transport Distance. We calculate the Wasserstein distance (Optimal Transport) [47] between the human and robot video embeddings, i.e., the cost of the optimal transport plan that transfers one sequence of video embeddings into another. The robot’s embedding distribution is defined as $\rho_{R}=\{1/T,1/T,\dots,1/T\}$, and the human’s embedding distribution is defined as $\rho_{H}=\{1/T^{\prime},1/T^{\prime},\dots,1/T^{\prime}\}$, where $T$ and $T^{\prime}$ are the lengths of the video sequences respectively. The cost function for the transport is $C\in\mathbb{R}^{T\times T^{\prime}}$ where $C^{ij}$ is the cosine distance between the robot embedding $z_{R}^{i}$ and the human embedding $z_{H}^{j}$. Our goal is to find the optimal assignment $M^{*}\in\mathbb{R}^{T\times T^{\prime}}$ that transports the distribution from $\rho_{R}$ to $\rho_{H}$ while minimizing the cost of the plan. Formally, we need to find $M^{*}=\operatorname*{arg\,min}\limits_{M}\sum_{i}\sum_{j}C^{ij}M^{ij}$. After solving the optimal transport assignment, the distance function is the cost of the plan, i.e., $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{i}\sum_{j}C^{ij}M^{*ij}$. In practice, we optimize an entropy-regularized version of this problem to find an approximate solution efficiently using the Sinkhorn-Knopp algorithm [47].
@@
Method 2: Temporal-Cyclic Consistency (TCC) Distance. We calculate the TCC loss between human and robot videos following  [48] which computes cycle consistency between robot video embeddings $\mathbf{z_{R}}=\{z_{R}^{1},z_{R}^{2},\dots,z_{R}^{T}\}$ and human video embeddings $\mathbf{z_{H}}=\{z_{H}^{1},z_{H}^{2},\dots,z_{H}^{T^{\prime}}\}$. For each robot frame $z_{R}^{t}$, we first compute a similarity distribution $\mathbf{\alpha}$ of $z^{R}_{t}$ with respect to the human’s embeddings, to find a soft-nearest neighbor $\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}$. Then, $\tilde{z}_{H}$ cycles back to the robot video by again computing its similarity distribution $\mathbf{\beta}$ with respect to robot video embeddings to get its soft-nearest neighbor $\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$. The TCC distance for a robot frame $z_{R}^{t}$ is the mean square error with its cycled-back frame $\tilde{z}_{R}^{t}$ as $\ell_{\rm tcc}=||z_{R}^{t}-\tilde{z}_{R}^{t}||_{2}$. We define the video-level TCC distance function by summing over the frame-level losses $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{t=1}^{T}\ell_{\rm tcc}(z_{R}^{t})$.
@@
We hypothesize that video retrieval using TCC distance can be inaccurate in two cases: (1) When human and robot embeddings differ due to variations in execution speed or style, leading to poor frame alignment. (2) When multiple robot embeddings correspond to a single human frame, for eg. unimanual robot tasks versus bimanual human actions.
@@ V Experiments
Figure 2: Performance on Mismatched Execution Datasets. We present results on three datasets (left). As the demonstrator’s execution deviates further from those of the robot, policies trained with our framework RHyME consistently outperforms XSkill measured by task recall and imprecision rates.
Figure 3: Realworld Results. (Left) Task Embeddings: We use t-SNE to visualize cross-embodiment latent embeddings from the human and robot completing three tasks. (Right) Task Completion: We compare the performance of RHyME with XSkill on seen and unseen long-horizon tasks specified by human prompt videos. Opaque segments indicate Task Completion rate, and augmented transparent bars indicate Task Attempt rate.
@@
Setup. Simulation: We evaluate our approach using the Franka Kitchen simulator [49], where a 7-DOF Franka arm performs $7$ different tasks. We generate $3$ cross-embodiment video datasets each progressively increasing the embodiment and execution mismatch, which contain $580$ long-horizon robot trajectories completing a sequence of 4 tasks and a bank of cross-embodiment demonstrator play data ($>$ 3 hours) for training our models. First, in Sphere-Easy, we replace the robot’s visual rendering with a sphere following the gripper’s position, creating a visual gap between robot and demonstrator. Second, in Sphere-Medium, we introduce manipulation style mismatches such as the robot dragging an object while the demonstrator lifts and carries it. Finally, in Sphere-Hard, we create a further divergence where the demonstrator performs two tasks simultaneously, similar to how humans use two hands. Fig. 2 (left) illustrates the cross-embodiment datasets. Realworld: We use a 7-DOF Franka arm to perform $4$ different tasks. We train our models on $40$ long-horizon robot trajectories completing a sequence of 3 tasks and $\sim$15 minutes of human play data, holding out unseen compositions of 3 test tasks.
@@
Baselines. XSkill [4] simply conditions on robot videos during train-time, and uses its shared representation space to zero-shot generalize to inputs of human videos at test-time. OraclePairing [7, 8] is the gold-standard approach, assuming an oracle pairs human demonstrations with robot trajectories, enabling conditioning on the human at train time. Our approach, RHyME, finds a middle ground. Without pairing, it imagines human videos that perform the same tasks as a robot trajectory (Section IV-B) by exploiting sequence-level correspondences. We compare two variants of our algorithm, RHyME-TCC and RHyME-OT, which differ in their distance functions used for retrieval. In the realworld, we use XSkill [4] as a baseline for RHyME-OT. We also show how vision representations can be improved using short-horizon robot-demonstrator task pairs.
@@
Evaluation and Success Metrics. At test-time, we provide the robot policy with long-horizon human videos as prompts. Simulation: We evaluate one-shot imitation performance across 20 different demonstrator videos, rolling out the robot policy from the last 5 model checkpoints, yielding 100 total trials per dataset. We measure Task Recall, which assesses recall by counting the successfully completed tasks shown in the demonstrator’s video, and Task Imprecision, which measures imprecision and reports the percentage of tasks the robot attempts incorrectly—those not specified in the demonstrator’s video. Realworld: We evaluate performance across 30 different human videos (20 seen, 10 unseen). We break down Task Recall into two metrics: Task Attempts and Task Completions, which measure (a) the robot’s ability to attempt tasks specified by the human video and (b) the low-level control policy’s ability to fully complete the tasks.
Figure 4: Cross-Embodiment Vision Embeddings. (Left) Visualizing task embeddings. We use t-SNE to visualize cross-embodiment latent embeddings generated by robot and demonstrator when executing different tasks on all three datasets. (Right) TCC Failure Example: The robot and video clip are equivalent, but specific frames have high TCC losses. For example, a frame showing the robot performing the ‘kettle’ action has a high loss due to its nearest neighbor in the video performing both ‘kettle’ and ‘light’ actions. This frame cycles back to the robot performing ‘light’, which is mismatched.
Q1. How does performance vary across different levels of execution mismatch? As the cross-embodiment demonstrator’s execution deviates further from those of the robot, policies trained with our framework RHyME consistently outperform XSkill (Fig. 2), with the largest gap in the bimanual demonstrator setting Sphere-Hard ($53\%$ vs $1\%$). The OraclePairing baseline serves as an upper bound on performance.
@@
Fig. 4 (left) investigates this trend by probing the visual representations of the video encoder $\phi$, common across policies. We plot the image embeddings of the robot and demonstrator across three different tasks using a t-SNE plot. As execution mismatch increases, the robot and demonstrator embeddings become less clustered by task, supporting XSkill’s inability to zero-shot transfer to demonstrator embeddings at test-time in Sphere-Hard. RHyME algorithms overcome this problem and successfully retrieve the correct demonstration videos at train-time by reasoning over sequences of embeddings. However, RHyME-OT outperforms RHyME-TCC across all three datasets in both metrics, suggesting inaccurate train-time retrievals with TCC.
Q2. How does RHyME perform on real kitchen tasks when prompted with human videos? With natural visual and execution mismatches between human and robot videos, RHyME consistently outperforms XSkill when prompted with both seen and unseen human prompt videos (Fig. 3 Right). We observe marginal benefits in Task Attempts and Task Completions when faced with the compositions of the tasks seen, but record significant improvements in both metrics (83% vs. 50% and 67% vs. 33%, respectively) in the unseen setting to which our framework aims to generalize.
@@
Figure 5: Optimal Transport Distances. We measure the similarity between robot and demonstrator videos on the Sphere-Hard dataset by computing the cost of the Optimal Transport (OT) plans. The sum over the entire transport cost matrix costs yields the distance between videos. OT costs are lowest when tasks are the same between videos (highlighted by a tick).
@@
Q3. How does video retrieval using Optimal Transport and TCC impact policies at test-time? As task embedding clusters deviate due to execution mismatches, we observe inaccuracies in TCC retrievals: in Sphere-Hard (Fig. 4 (right)) when both clips complete the same two tasks, a bimanual task embedding lies in between two robot task clusters which results in cycling-back to the incorrect robot frame, leading to high task imprecision. RHyME-OT performs strictly better across datasets (Fig. 2). The key reason for this performance difference is that optimal transport computes distances by matching videos across a sequence of embeddings. Fig. 5 visualizes the cost of the optimal transport plan between prompt robot clips and demonstration videos in the hard Sphere-Hard dataset. Comparing a robot clip doing two tasks (e.g. kettle and light), the transport cost across assignments is minimum only when compared to the demonstrator performing those same two tasks. TCC, on the other hand, attempts to establish one-to-one correspondences between the robot and demonstration frames, which are lacking in this dataset.
Q4. Where does RHyME succeed, and what are common failure modes of other methods? We visualize the vision embeddings using t-SNE (Fig. 3 Left). We find that image embeddings in the realworld are generally clustered by task, but tend to deviate during the Light Switch task. Consequently, we observed that XSkill never attempts the task when prompted with human embeddings. On the other hand, in unseen settings, RHyME always attempts the Light Switch task and completed it 9 out of 10 times. The Optimal Transport retrieval (Sec. IV) used to imagine the paired dataset recognizes can correctly match human and robot clips completing the same task by reasoning over the distribution of embeddings rather than relying on perfect embedding alignment, so RHyME is able to accurately pair action labels with imagined human videos at train time and obtain better performance (Fig. 3 Right).
@@
Q5. Does fine-tuning visual representations with task-equivalent pairs improve one-shot imitation? Following Section IV-A, we assume access to short-horizon task pairings across embodiments and apply $L_{task}(\phi)$ on vision representations in the Sphere-Hard setting. We find that encouraging induced distributions over embeddings to be similar lifts the performance of both XSkill and RHyME-OT (Fig. 6), and scales up with more paired clips. Ultimately, comparing induced distributions over embeddings with optimal transport is a beneficial design choice for matching clips at a task-level in the face of execution mismatches, as RHyME-OT (0% fine-tuned) still significantly outperforms XSkill (fine-tuned).